# Fast, robust and laser-free universal entangling gates for trapped-ion quantum computing

### v1

@@ III.2 Coupling to the motional degrees-of-freedom
We see in Fig. 3 the recovery of the Ramsey fringe contrast after the full duration of the entangling gate’s time ($\approx 300\mu$s). At this time, the ion’s internal qubit states are disentangled again from the vibrational motion. This is the desired condition after an entangling gate has been completed, as described in Sec. IV. For comparison, at $\Omega_{1}=2\pi\times 61\leavevmode\nobreak$kHz where the ion-phonon coupling is weak, a smaller reduction of the fringe contrast in Fig. 3 occurs.
@@
Figure 3: Entanglement and disentanglement of qubit states and motional states of a single ion by application of a phase modulated field (Eq.(1)) for the duration of a two-qubit gate. For Rabi frequency $\Omega_{1}=2\pi\times 94.8$kHz, close to the secular trap frequency, entanglement is signified by a maximal reduction of Ramsey fringe contrast at $t=150\mu$s, which corresponds to approximately half the two-qubit gate time. At $t=300\mu$s, qubit and motion are disentangled again, signified by a recovery of the fringe contrast. For Rabi frequency $\Omega_{1}=2\pi\times 61.0$kHz, qubit-phonon coupling, and, therefore, qubit-phonon entanglement is smaller, resulting in a smaller reduction of fringe contrast halfway through the gate. The dotted lines are added to guide the eye.
@@ IV Two-Qubit entangling gate
IV.1 Experimental procedure
The qubits are first optically pumped into the state $|00\rangle$. Subsequently, they are both initialized in the super-position state $\ket{++}$ by applying a resonant $\pi/2$-pulse to each of the qubits. Then, the phase-modulated driving field (Eq. (1)) is applied to both qubits for duration $t$ with Rabi frequency $\Omega_{1}$ set close to the COM mode frequency, $\nu$. The difference $\epsilon$ between $\nu$ and $\Omega_{1}$ is determined by the effective Lamb-Dicke parameter, $\eta$ and by $\nu$ as detailed in Sec. IV.3, Eq. (10). Here, we use $\Omega_{1}=2\pi\times 94.8$ kHz to fulfill this condition. $\Omega_{2}$, characterizing the phase modulation depth, is set to $\Omega_{2}=2\pi\times 71$ kHz, for double dressing and protecting the qubit’s coherence against RF amplitude fluctuations. For a chosen $\Omega_{1}$, the value of $\Omega_{2}$ is given by assumptions described in Sec. IV.3 and guided by numerical simulations, searching for the minimal residual excitation of the ions’ motion. To reconstruct the density matrix $\rho$, after application of the gate, tomography of the two- qubit state is carried out. For this purpose, the sequence of initial state preparation, subsequent two-qubit entangling gate, and measurement is repeated $M$ times $(200<M<600)$. After preparation of the entangled state, in order to effectively measure in different bare state bases, different sets of local single-qubit rotations are applied before a projective measurement in the $z-$basis takes place [26] (for details see appendix C).
IV.2 Experimental results
A full reconstruction of the two qubit density matrix allows calculating the purity $\mathcal{P}=\text{tr}(\rho^{2})$ as well as the negativity $\mathcal{N}$ of the quantum state as basis-independent measures. Here, the purity of a state indicates the entanglement between the qubits and their state of motion. For qubit and motional states that are perfectly disentangled at the end of the gate as desired, the purity of the state is unity, while deviations from unity are attributed to residual entanglement with phonons. For a maximally entangled state (such as a Bell state $|\Phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ and $|\Psi^{+}\rangle=(|01\rangle+|10\rangle)/\sqrt{2}$) the negativity is 0.5. The evolution of negativity from zero (the unentangled 2-qubit state) towards the value of 0.5 (that indicates a maximally entangled state) reaching the experimental maximum of $0.48^{+2}_{-6}$, is shown in Fig. 4.
The statistical error of the negativity is calculated as standard error of a set of numerically sampled density matrices. Here, we assume the entries of the reconstructed density matrix to be the mean value of a normal distribution with the width of the standard error assigned during the reconstruction (see appendix D for details). Due to this reconstruction, the error interval might exceed physical limits. In this case, we truncate the statistical error bars of physical quantities. Similarly, we truncate the error interval of fidelity and purity values which are based on Gaussian propagation of normal distributions.
An oscillation in the state’s purity between $\approx 0.6$ and 1.0 is evident in Fig. 4. This is due to varying qubit-phonon entanglement during the gate evolution time. This oscillation is a direct indication of the phase space trajectory generated by the gate, where at half the gate time, the purity is minimal and at the gate time, the purity is recovered (compare also Fig. 3). This indicates that one closed loop in phase space during the evolution of the gate has been completed, realizing the fastest possible two-qubit gate for the parameter set used here.
Figure 4: Time evolution of the two-qubit entangling gate quantified by the two-qubit state’s negativity and purity, calculated from a full state reconstruction at interaction times $t$. Throughout the gate evolution, the negativity increases to $0.48^{+2}_{-6}$ at time $t=313\leavevmode\nobreak\ \mu$s. This evolution time creates the Bell state explicitly shown in Fig. 5. The oscillation between purity of $\approx 0.6$ and 1.0 indicates varying qubit-phonon entanglement during the gate evolution signifying no entanglement between qubits and motional states at the beginning and end of the gate sequence. Here the achieved pururity at the gate time $t=313\leavevmode\nobreak\ \mu$s is $0.99^{+1}_{-3}$ Based on Eq. (1), the gate evolution is simulated using $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, $\Omega_{2}=2\pi\times 71\leavevmode\nobreak$kHz, and a trap frequency of $\nu=2\pi\times 98.08$ kHz yielding the gate time $t_{\text{g}}^{\text{sim}}=309.5\leavevmode\nobreak\ \mu$s.
Based on the time evolution measurements of the product-state excitation probability in the bare state basis (see appendix E), we select a gate time where the product state excitation probability matches the Bell state $|\Phi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0.5$ and $P_{|01\rangle}=P_{|10\rangle}=0$, or the Bell state $|\Psi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0$ and $P_{|01\rangle}=P_{|10\rangle}=0.5$. The gate time for generating a the Bell state $|\Phi^{+}\rangle$ is $t^{\mathrm{exp}}_{\mathrm{g}}=$ 313 $\mathrm{\mathrm{\mu}s}$, while the Bell state $|\Psi^{+}\rangle$ is generated after time ($t^{\mathrm{exp}}_{\mathrm{g}}=$ 310 $\mathrm{\mathrm{\mu}s}$). Fig. 5 shows the reconstructed density matrices, corrected for state detection errors (details are given in appendix B).
For $t^{\mathrm{exp}}_{g}=313\leavevmode\nobreak\ \mathrm{\mu}$s, we report measurement outcomes of $\mathcal{N}=0.48^{+2}_{-6}$ and $\mathcal{P}=0.99^{+1}_{-3}$. Once the density matrix is reconstructed, numerically an optimal set of single qubit rotations is computed to rotate the entangled state into a desired Bell state. While the entangling gate’s parameters are chosen such that the produced state is close to |$\ket{\psi}=\ket{\Phi^{+}}(\ket{\Psi^{+}})$, these numerical single-qubit rotations are applied to infer the resulting state fidelity as
where $R_{j}(\alpha_{j},\beta_{j},\gamma_{j})$ represents the rotation of qubit $j$ around the angles $\alpha_{j},\beta_{j},$ and $\gamma_{j}$. A fidelity of $\mathcal{F}=98^{+2}_{-3}$ w.r.t a maximally entangled $|\Phi^{+}\rangle$ Bell state is achieved (see appendix D). For a gate time of $t^{\mathrm{exp}}_{g}=310\leavevmode\nobreak\ \mathrm{\mu}$s, we report $\mathcal{N}=0.47^{+3}_{-6}$ and $\mathcal{P}=0.98^{+2}_{-3}$, and $\mathcal{F}=97^{+3}_{-3}$ w.r.t a maximally entangled $|\Psi^{+}\rangle$ Bell state. Both Bell states can be selectively generated by choosing appropriate gate times.
IV.3 Gate description
The idea to use double-dressed states for two-qubit gates was first proposed in [27]. Our gate extends that scheme using the phase modulation as an effective second drive, making it robust to amplitude fluctuations. The gate operation can be readily understood in terms of its optical counterparts [1, 28, 29]. Let us observe Eq. (3) with $\Omega_{2}=0$. The Hamiltonian is similar to the Cirac-Zoller interaction Hamiltonian [1], where $\Omega_{1}$ replaces the detuning of the laser drive frequency from the optical transition (e.g. qubit) frequency and where the carrier transition is eliminated. Therefore, in an analogous way, we can implement an entangling interaction by choosing $\Omega_{1}$ to be close to the motional sidebands. The effective second drive is then used to decouple the motional mode dependent frequency shift turning the interaction into a Mølmer-Sørensen-type gate.
@@ IV.3 Gate description
where the stretch mode is neglected as the coupling to stretch mode is smaller, and it is far detuned.
Both ions are driven with a phase-modulated drive described by the Hamiltonian given in Eq. (1), where we set $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}=\Omega_{1}$ for both qubits. Following similar transformations as for the one qubit interaction (see appendix A), we arrive at the double dressed state Hamiltonian
@@
Here, we define the double-dressed basis operators $F_{z}=S_{x},\ F_{y}=S_{y},\ F_{z}=-S_{z}$, and we assume the rotating wave approximation (RWA) like in the single-qubit interaction. Denoting $\Omega_{1}=\nu-\epsilon$, the term proportional to $F_{y}^{(j)}$ in Eq. (6) is negligible, assuming that $|\epsilon\pm\frac{\Omega_{2}}{2}|\gg\frac{\eta\nu}{2}$, $|\Omega_{1}+\nu-\frac{\Omega_{2}}{2}|\gg\frac{\eta\nu}{2}$, and $|\epsilon|\ll\Omega_{1}+\nu$. The interaction is simplified to
@@
which is a Mølmer-Sørensen-like Hamiltonian, that creates the effective interaction
@@
For the present axial trap frequency and static magnetic gradient, this analytical gate time of 309.8 $\mu$s matches well the experimental observed one. With the simulated gate time of 309.5 $\mu$s, corresponding to the time when phonon entanglement is minimized, we expect a unitary gate fidelity of 99.3 %. The residual infidelity is attributed to the quantum stark shift remaining in the double dressed basis, which is on the order of $(\eta\nu)^{2}/4\Omega_{2}$. Therefore, the infidelity scales as $(\pi\eta\nu/2\Omega_{2})^{2}\sim 0.7\%$ By parameter optimization or pulse shaping methods this can be further reduced. The remaining infidelity in the measured entangling gates is explained by the effect of phonon heating.

### v2

@@ III.2 Coherent coupling to motional degrees-of-freedom
The recovery of the Ramsey fringe contrast after $\approx 300\leavevmode\nobreak\ \mu$s shows that the ion’s internal qubit states are disentangled again from the vibrational motion. This is the desired condition after an entangling gate has been completed (details in Sec. IV). For comparison, at $\Omega_{1}=2\pi\times 61\leavevmode\nobreak$kHz where the ion-phonon coupling is weak, a smaller reduction of the fringe contrast occurs (red data points in Fig. 3). For this setting, the qubit-phonon entanglement is reduced as compared to the case when the Rabi frequency $\Omega_{1}$ (which determines the energy splitting between the single dressed states) is set close to the motional frequency (blue data points).
@@
Figure 3: Entanglement and disentanglement of qubit states and motional states of a single ion by application of a phase modulated field (Eq.(2)). For Rabi frequency $\Omega_{1}=2\pi\times 94.8$kHz, close to the secular trap frequency, entanglement is signified by a significant reduction of Ramsey fringe contrast at $t=150\mu$s (blue data points). At $t=300\mu$s, qubit and motion are disentangled again, signified by a recovery of the fringe contrast. For Rabi frequency $\Omega_{1}=2\pi\times 61.0$kHz, qubit-phonon coupling, and, therefore, qubit-phonon entanglement is smaller, resulting in a smaller reduction of fringe contrast (red data points). The dotted lines are added to guide the eye.
@@ IV Two-qubit entangling gate
We now apply two phase-modulated fields, each of which is tuned to the resonance of one of the two ionic qubits.
IV.1 Gate description
The idea to use double-dressed states for two-qubit gates was first proposed in [57]. Our gate extends that scheme using phase modulation of a single driving field as an effective second drive, making it robust to amplitude fluctuations.
@@ IV.1 Gate description
where the stretch mode is neglected as the coupling to this mode is far detuned.
Both ions are driven with a phase-modulated drive such that the drive Hamiltonian reads
$H_{D}^{(2)}=\sum_{j=1,2}$ $\Omega_{1}^{(j,\text{Amp})}\sigma^{(j)}_{x}$ (5)
×$\times\cos\left(\omega_{0}^{(j)}t+\frac{\Omega_{2}}{\Omega_{1}^{(j,\text{Phase})}}\sin\left(\Omega_{1}^{(j,\text{Phase})}t\right)\right).$
Here, $\Omega_{1}^{(j,\text{Phase})}$ is the phase modulation frequency of the $j$-th qubit and $\Omega_{1}^{(j,\text{Amp})}$ is the Rabi frequency determined by the RF field amplitude for qubit $j$. All frequencies are angular frequencies.
We set $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}=\Omega_{1}$ for both qubits. Then, following similar transformations as for the one qubit interaction (appendix A), we arrive at the double dressed state Hamiltonian
@@
Here, we define the double-dressed basis operators $F_{z}=S_{x}=-\sigma_{z},\ F_{y}=S_{y}=\sigma_{y},\ F_{x}=-S_{z}=-\sigma_{x}$, and we assume the rotating wave approximation (RWA) like in the single-qubit interaction.
We denote the difference between motional mode frequency and Rabi frequency by $\epsilon=\nu-\Omega_{1}$. The term proportional to $F_{y}^{(j)}$ in Eq. (6) may be interpreted as a driving term in the double-dressed basis, driving sidebands with frequency $\Omega_{1}$. Assuming that
$|\epsilon\pm\frac{\Omega_{2}}{2}|\gg\frac{\eta\nu}{2},\\ |\Omega_{1}+\nu-\frac{\Omega_{2}}{2}|\gg\frac{\eta\nu}{2},\\ \text{and}|\epsilon|\ll\Omega_{1}+\nu,$ (7)
this driving field is far-detuned from the sideband’s transition frequencies at $\Omega_{2}/2\pm\nu$ and the term proportional to $F_{y}$ is negligible. The interaction then simplifies to
@@
which is a Mølmer-Sørensen-type Hamiltonian, that creates the effective interaction
@@
In this case, the effective interaction is given by $-\frac{\eta\nu}{8}F_{z}^{(1)}F_{z}^{(2)}$, that is, the effective coupling between the two qubits is directly proportional to the Lamb-Dicke parameter and, therefore, proportional to the static magnetic gradient.
The conditional gate operation described above can be interpreted in terms of its optical counterparts [1, 47, 48, 46]. Let us observe Eq. (3) with $\Omega_{2}=0$. The Hamiltonian is then similar to the Cirac-Zoller interaction Hamiltonian [1], where $\Omega_{1}$ replaces the detuning of the laser drive frequency from the optical qubit resonance and where the carrier transition is eliminated. Therefore, in an analogous way, we can implement an entangling interaction by choosing $\Omega_{1}$ to be close to the motional sidebands. The effective second drive is then used to decouple the motional mode dependent frequency shift turning the interaction into a Mølmer-Sørensen (MS)-type gate.
IV.2 Experimental procedure
In order to realize this gate, the qubits are first optically pumped into state $|00\rangle$. Subsequently, they are both initialized in the super-position state $\ket{++}$ by applying a resonant $\pi/2$-pulse to each of the qubits. Then, the phase-modulated driving fields (Eq. (5)) are applied to both qubits.
For our gate scheme, we require $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}\equiv\Omega_{1},$ (optimization over the later condition may result in a longer coherence time [56]). To ensure $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}$, RF-amplitudes are calibrated by recording and fitting Rabi oscillations of each individually addressed qubit (appendix C).
The phase modulated field is applied for duration $t$ with the Rabi frequency $\Omega_{1}$ set close to the COM mode frequency, $\nu$: The detuning $\epsilon$ between $\nu$ and $\Omega_{1}$ is set such as to satisfy condition Eq. (11). Here, we use $\Omega_{1}=2\pi\times 94.8$ kHz to fulfill this condition. $\Omega_{2}$, characterizing the phase modulation depth, is set to $\Omega_{2}=2\pi\times 71$ kHz, for double dressing and protecting the qubit’s coherence against RF amplitude fluctuations. For a chosen $\Omega_{1}$, the considerations for choosing $\Omega_{2}$ are described in Sec. IV.1 and guided by numerical simulations, searching for the minimal residual excitation of the ions’ motion.
To reconstruct the density matrix $\rho$, after application of the gate, tomography of the two-qubit state is carried out. For this purpose, the sequence of initial state preparation, subsequent two-qubit entangling gate, and measurement is repeated $M$ times $(200<M<600)$. After preparation of the entangled state, in order to effectively measure in different bare state bases, different sets of local single-qubit rotations are applied before a projective measurement in the $z-$basis takes place [58] (appendix D).
IV.3 Experimental results
A full reconstruction of the two qubit density matrix allows calculating the purity $\mathcal{P}=\text{tr}(\rho^{2})$ as well as the negativity $\mathcal{N}$ of the quantum state as basis-independent measures [59]. Here, the purity of a state is a measure for the degree of residual entanglement between the qubits’ internal atomic state and their external motional state. For qubit and motional states that are perfectly disentangled at the gate time as desired, the purity of the state is unity, while deviations from unity are attributed to residual entanglement with phonons. For a maximally entangled state (such as a Bell state $|\Phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ and $|\Psi^{+}\rangle=(|01\rangle+|10\rangle)/\sqrt{2}$) the negativity is 0.5. The evolution of negativity from zero (the unentangled 2-qubit state) towards the value of 0.5 (that indicates a maximally entangled state) reaching the experimental maximum of $0.48^{+2}_{-6}$, is shown in Fig. 4.
The statistical error of the negativity is calculated as the Gaussian standard deviation of a set of numerically sampled density matrices. Here, we assume that the entries of the reconstructed density matrix are the mean value of a normal distribution having the width of the standard deviation assigned during the reconstruction (appendix E). Due to this reconstruction, the error interval might exceed physical limits. In this case, we truncate the statistical error bars of physical quantities. Similarly, we truncate the error interval of fidelity and purity values based on Gaussian propagation of normal distributions.
An oscillation in the state’s purity from initially 1.0 to $\approx$ 0.6 and back to 1.0 is evident in Fig. 4. This is due to varying qubit-phonon entanglement during the gate evolution time. This oscillation is a direct indication of the phase space trajectory generated by the gate, where at half the gate time, the purity is minimal and at the gate time, the purity is recovered (compare also Fig. 3). This indicates that one closed loop in phase space during the evolution of the gate has been completed [48, 50], realizing the fastest possible two-qubit gate for the parameter set used here (compare section IV.1, Eq. 10, where different gate times are discussed).
Figure 4: Time evolution of the two-qubit entangling gate quantified by the two-qubit state’s negativity (blue data points) and purity (red data points), calculated from a full state reconstruction at interaction times $t$. Here, the negativity increases to $0.48^{+2}_{-6}$ at time $t=313\leavevmode\nobreak\ \mu$s. At this evolution time the Bell state $|\Psi^{+}\rangle$ is created whose experimentally reconstructed density matrix is shown in Fig. 5. The oscillation between a purity of $\approx 0.6$ and 1.0 indicates varying qubit-phonon entanglement during the gate evolution signifying no entanglement between qubits and motional states at the beginning and end of the gate sequence. Here, the achieved purity at the gate time $t=313\leavevmode\nobreak\ \mu$s is $0.99^{+1}_{-3}$. Based on Eq. (5), the gate evolution is simulated using $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, $\Omega_{2}=2\pi\times 71\leavevmode\nobreak$kHz, and a trap frequency of $\nu=2\pi\times 98.08$ kHz. The solid lines show the simulated gate evolution with gate time $t_{\text{g}}^{\text{sim}}=309.5\leavevmode\nobreak\ \mu$s. Discrepancies between experimental data and simulation are caused by fluctuations in experimental parameters (notably the not stabilized secular trap frequency) on a daily basis.
Based on the time evolution measurements of the product-state excitation probability in the bare state basis (appendix F), we select a gate time where the product state excitation probability matches the Bell state $|\Phi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0.5$ and $P_{|01\rangle}=P_{|10\rangle}=0$, or the Bell state $|\Psi^{+}\rangle$, i.e. $P_{|00\rangle}=P_{|11\rangle}=0$ and $P_{|01\rangle}=P_{|10\rangle}=0.5$.
The gate time for generating the Bell state $|\Phi^{+}\rangle$ is $t^{\mathrm{exp}}_{\mathrm{g}}=313\leavevmode\nobreak\ \mathrm{\mathrm{\mu}s}$, while the Bell state $|\Psi^{+}\rangle$ is generated after time ($t^{\mathrm{exp}}_{\mathrm{g}}=310\leavevmode\nobreak\ \mathrm{\mathrm{\mu}s}$). Fig. 5 shows reconstructed density matrices for these Bell states, corrected for state detection errors (appendix G). For $t^{\mathrm{exp}}_{g}=313\leavevmode\nobreak\ \mathrm{\mu}$s, we report measurement outcomes of $\mathcal{N}=0.48^{+2}_{-6}$ and $\mathcal{P}=0.99^{+1}_{-3}$.
Once the density matrix is reconstructed, numerically an optimal set of single qubit rotations is computed to rotate the entangled state into a desired Bell state. While the entangling gate’s parameters are chosen such that the produced state is close to |$\ket{\psi}=\ket{\Phi^{+}}(\ket{\Psi^{+}})$, these numerical single-qubit rotations are applied to infer the resulting state fidelity as
where $R_{j}(\alpha_{j},\beta_{j},\gamma_{j})$ represents the rotation of qubit $j$ around the angles $\alpha_{j},\beta_{j},$ and $\gamma_{j}$. A fidelity of $\mathcal{F}=98^{+2}_{-3}$ w.r.t a maximally entangled $|\Phi^{+}\rangle$ Bell state is achieved. For a gate time of $t^{\mathrm{exp}}_{g}=310\leavevmode\nobreak\ \mathrm{\mu}$s, we report $\mathcal{N}=0.47^{+3}_{-6}$ and $\mathcal{P}=0.98^{+2}_{-3}$, and $\mathcal{F}=97^{+3}_{-3}$ w.r.t a maximally entangled $|\Psi^{+}\rangle$ Bell state. Both Bell states can be selectively generated by choosing appropriate gate times.
@@ IV.3 Experimental results
The calculated gate time of 309.8 µs, obtained using the aforementioned experimental parameters and the analytical expression given in Eq. 11, is in close agreement with the experimental result. By simulating the system and driving Hamiltonian (Eq. 4 and Eq. 5) after applying only the RWA w.r.t the bare state energies $H_{0}=\sum_{i}\omega_{0}^{i}/2\sigma_{z}^{i}$, we yield the numerically gate time of 309.5 $\mu$s, corresponding to the time when phonon entanglement is minimized. From these simulations we expect a unitary gate fidelity of 99.3 %. The residual infidelity is attributed to the quantum stark shift remaining in the double dressed basis, which is on the order of $(\eta\nu)^{2}/4\Omega_{2}$. Therefore, the infidelity scales as $(\pi\eta\nu/2\Omega_{2})^{2}\sim 0.7\%$ By parameter optimization or pulse shaping methods this can be further reduced. The remaining infidelity in the measured entangling gates is explained by the effect of phonon heating.
IV.4 Robustness
We investigate the robustness of the double dressed entangling gate against symmetric (asymmetric) detunings $\delta_{1}$ ($\delta_{2}$) of $\Omega^{\text{Amp}}_{1}$ w.r.t. $\nu-\epsilon$ by measuring the entangled state negativity at a gate time of 313 $\mathrm{\mu}$s with $\Omega_{1}^{\text{Phase}}=2\pi\times 94.8$ kHz. Fig. 6(a) shows the effect of a mismatch $\delta_{1}$ between the physical Rabi frequency $\Omega_{1}^{\text{Amp}}$ for both ions and the phase modulation frequency $\Omega_{1}^{\text{Phase}}$ ($\delta_{1}=\Omega_{1}^{\text{Phase}}-\Omega_{1}^{\text{Amp}}$). Fig. 6(b) shows the impact of a mismatch of the physical Rabi frequencies between the two ions ($\delta_{2}=\Omega_{1}^{\text{Amp},1}-\Omega_{1}^{\text{Amp},2}$), while $\Omega_{1}^{\text{Amp},1}=\Omega_{1}^{\text{Phase}}$ holds.
For $\delta_{1}\approx 3\leavevmode\nobreak\ \%$ and $\delta_{2}\approx 8\leavevmode\nobreak\ \%$ of $\Omega_{1}^{\text{Phase}}$, the measured negativity stays above $80\leavevmode\nobreak\ \%$ of the optimal negativity, showing the robustness of the entangling gate against such mismatches of the RF field amplitude. Similar results were observed for other RF driven MS-type gates [60], quantified in terms of fidelity. For small detunings, as typically occur in experiments, a quadratic dependence between negativity and detuning is observed, showing the stability against small detunings. The robustness of the gate, made it possible to scan a wide range of $\delta_{1,2}$, exceeding typical uncertainties during the experiments by two orders of magnitude.